Let f(x) = x + 2 and g(x) = x2 Find ( g o f)(-5)

Answer by jim_thompson5910(35256)

(g o f)(-5) means g(f(-5)). So we need to find g(f(x)) first. Then we can plug in x = -5

g(x) = x^2

g(f(x)) = (f(x))^2

g(f(x)) = (x+2)^2

(g o f)(x) = (x+2)^2

(g o f)(-5) = (-5+2)^2

(g o f)(-5) = (3)^2

(g o f)(-5) = 9

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim